Qd(=50-8P) and Qs(=17.5+10P)


1
Expert's answer
2022-04-25T16:58:03-0400

"Q_d=50-8P,"

"Q_s=17.5+10P,"

"Q_d=Q_s,"

"50-8P=17.5+10P,"

"32.5=18P,"

"P=1.8,"

"Q_d=Q_s=35.6."